# FARM: Functional Group-Aware Representations for Small Molecules

### v2

@@ Abstract
We introduce Functional Group-Aware Representations for Small Molecules (FARM), a novel foundation model designed to bridge the gap between SMILES, natural language, and molecular graphs. The key innovation of FARM lies in its functional group-aware tokenization, which directly incorporates functional group information into the representations. This strategic reduction in tokenization granularity is intentionally aligned with key drivers of functional properties (i.e., functional groups), enhancing the model’s understanding of chemical language. By expanding the chemical lexicon, FARM more effectively bridges SMILES and natural language, ultimately advancing the model’s capacity to predict molecular properties. FARM also represents molecules from two perspectives: by using masked language modeling to capture atom-level features and by employing graph neural networks to encode the whole molecule topology. By leveraging contrastive learning, FARM aligns these two views of representations into a unified molecular embedding. We rigorously evaluate FARM on the MoleculeNet dataset, where it achieves state-of-the-art performance on 10 out of 12 tasks. These results highlight FARM’s potential to improve molecular representation learning, with promising applications in drug discovery and pharmaceutical research.
@@ 1 Introduction
Artificial intelligence (AI) has emerged as a transformative tool in accelerating scientific discovery, particularly in drug development. It is increasingly employed for tasks such as molecular property prediction, drug-target interaction prediction, and quantitative structure-activity relationship (QSAR) modeling (Chen et al., 2016; Wen et al., 2017; Shen & Nicolaou, 2019; Walters & Barzilay, 2020; Achary, 2020; Wang et al., 2022a; Edwards et al., 2022; Zhang et al., 2023a; Nguyen et al., 2024a; Edwards et al., 2024b; a). However, one of the central challenges in this field is the scarcity of large labeled datasets required for traditional supervised learning methods. This has shifted the focus towards self-supervised pre-trained models that can extract meaningful patterns from vast amounts of unlabeled molecular data (Shen & Nicolaou, 2019). As a result, the development of robust foundation models for molecular representations is now more critical than ever. Despite significant advancements in other domains, such as natural language processing (NLP) and computer vision, there is still no dominant foundation model tailored to molecular representation in drug discovery (Zhang et al., 2023b). This paper begins to address this pressing gap by introducing an innovative approach that leverages functional group (FG)-aware tokenization in the context of both sequence-based and graph-based molecular representations. Analogous to what has been seen with prior studies, aligning these two representations generates a unified and comprehensive molecular embedding that effectively captures both atom-level features and the structural topology of molecules. The unique innovation here is that the reduction in tokenization granularity in a way that is intentionally interfaced with key drivers of functional properties (i.e., functional groups) enables this form of molecular embedding to better promote models’ capacity to understand and predict molecular functions. Figure 1 shows an overview of our molecular representation learning model.
Figure 1: Overview of FARM’s molecular representation learning framework.
Molecular structures are commonly represented either as sequences, like SMILES or SELFIES, or as molecular graphs. However, relying solely on one type of representation—whether sequence-based or graph-based—limits the ability to capture the full complexity of molecular structures. Sequence representations like SMILES can leverage powerful language models, such as BERT (Devlin, 2018) and GPT (Radford, 2018), which have proven highly successful in NLP tasks due to their ability to capture complex patterns, contextual relationships, and semantic nuances in sequential data. However, SMILES strings inherently lose crucial spatial and topological information, such as stereochemistry, 3D coordinates, and chirality—features that are often critical for accurate molecular predictions. On the other hand, graph neural networks (GNNs) excel at capturing the local topological structure of molecules but struggle with capturing long-range dependencies, such as interactions between distant atoms (Xu et al., 2018).
Our approach overcomes these limitations by integrating the strengths of both representations. We employ masked language models (MLMs) to capture robust atom-level features from SMILES while simultaneously using GNNs to model the structural topology of the molecule. These two representations are aligned through contrastive learning, resulting in a molecular embedding that comprehensively captures both atom-level and structural information. This alignment enables our model to fully represent molecular intricacies, leading to significant improvements in performance across a range of downstream tasks in cheminformatics.
The key novelty of this work is the integration of functional group information into molecular representations. We use the term functional groups to unify and expand upon related concepts such as “functional groups,” “motifs,” “fragments,” “substructures,” and “building blocks.” A functional group refers to a chemically meaningful portion of a molecule that contributes to its properties and behavior. This can include simple functional groups like hydroxyl (-OH) as well as more complex molecular substructures, such as ring systems, that serve distinct functional roles within the molecule. Functional groups (FGs) play a crucial role in determining a molecule’s properties, as illustrated in Figure 2. It presents the example of salicylic acid and aspirin, two molecules that share the same core structure but differ by just one functional group. This minor modification has profound implications, leading to large differences in their chemical properties and biological activities.
We introduce FG-aware tokenization and fragmentation, a fine-grained method that enhances molecular representations by incorporating detailed functional group information. This technique applies to both sequence and graph-based models, enriching the molecular representation with chemically meaningful context. The FG-aware tokenization and fragmentation assign each atom to its corresponding functional group, embedding chemical semantics directly into the molecular representation. This approach addresses a major limitation of sequence-based models, which typically focus only on individual atom types while neglecting the higher-level functional groups crucial for accurate molecular understanding. FG-aware tokenization enriches the SMILES representation with chemically relevant context, bridging the gap between the expansive vocabularies used in natural language models and the limited chemical lexicon typically available in molecular models.
Negative transfer occurs when a model trained on one domain or task struggles to generalize effectively to another due to differences in the underlying patterns. In molecular modeling, using only individual atom tokens without functional group information can lead to poor learning because the model fails to capture the full context of molecular behavior. The small, simplistic vocabulary of atom types, typically around 100 tokens, limits the model’s ability to understand complex chemical properties, often resulting in incorrect generalizations or degraded performance across different molecules and tasks. To address this, our FG-aware tokenization expands the vocabulary from 93 tokens to approximately 14,741 tokens by incorporating functional group information. While this significantly increases the complexity of the model, making training slower and harder to converge, it also prevents negative transfer by enabling the model to learn richer and more meaningful chemical semantics. This larger, more nuanced vocabulary allows the model to better capture the functional roles of atoms within molecules, improving its ability to generalize across tasks and ultimately leading to more accurate molecular representations.
Figure 2: Example of a pair of molecules, salicylic acid and aspirin, that share the same core structure but differ in a single functional group—where the hydroxyl group (-OH) in salicylic acid is replaced by an ester group (-COO-) in aspirin. This small change leads to significant differences in their chemical properties and biological activity.
To further advance the molecular representation, we focus on learning the structural aspects of the molecules. We utilize a FG knowledge graph to capture effective embeddings for each functional group, which are then used to learn structural representations of the molecule through link prediction between functional groups. This process ensures that the structural relationships among different functional groups are accurately captured and integrated into the final representation.
By aligning sequence-based and graph-based representations through contrastive loss, our approach achieves state-of-the-art results on 10 out of 12 benchmark tasks in the MoleculeNet dataset (Wu et al., 2018), demonstrating its robustness and versatility. These results underscore the potential of our pre-trained foundation model to significantly advance molecular representation learning, providing a powerful tool for addressing complex challenges in drug discovery and cheminformatics.
@@
2 Related Work
2.1 Functional group-Aware Molecular Representations
In recent years, there has been growing recognition that incorporating functional group (FG) information into molecular representations can significantly enhance model performance in downstream tasks. Approaches in this domain can be broadly classified into two categories: those leveraging language models (LMs) (Li et al., 2023; Xia et al., 2023) and those utilizing graph neural networks (GNNs) (Zhang et al., 2020; 2021; Yu & Gao, 2022; Yang et al., 2022; Wang et al., 2023; Wu et al., 2023; Han et al., 2023; Fang et al., 2023). Within these categories, methods either employ rule-based functional group detection, relying on predefined chemical rules to identify FGs, or adopt unsupervised strategies that infer substructures or motifs from the data. Regardless of the approach, these models consistently demonstrate that enriching molecular representations with functional group information leads to improved performance across a wide range of molecular property prediction tasks (Fang et al., 2023; Han et al., 2023; Wang et al., 2023). This underlines the importance of functional group awareness in achieving accurate and generalizable molecular representations.
In studies such as Zhang et al. (2021); Han et al. (2023); Chen et al. (2024); Yang et al. (2022), the BRICS algorithm (Degen et al., 2008) is employed to fragment molecules, with some extending the approach through additional rules to achieve finer-grained segmentation. However, BRICS relies on 16 predefined bond-breaking rules rooted in retrosynthetic reactions, which, while useful for synthesis planning, often result in coarse-grained fragmentation. This method focuses on general reaction-based bond breaking, rather than targeting functional group-specific characteristics, and therefore may overlook the detailed structural nuances needed for more precise molecular analysis. Additionally, while BRICS groups atoms into fragments, it does not explicitly identify or label functional groups, limiting its utility in tasks that require functional groups-specific information.
Other works (Li et al., 2023; Chen et al., 2024) leverage RDKit (Landrum, 2010) for functional group detection. While RDKit is effective in identifying common functional groups, its capabilities are limited to well-known groups and do not extend to detecting more complex functional groups, such as ring systems that often dominate molecular datasets and play a critical role in molecular function. Hence, RDKit may overlook less frequent functional groups or larger, more intricate structures that are essential for achieving a comprehensive molecular representation in diverse chemical datasets.
In Wang et al. (2023), the authors propose an unsupervised motif-based graph representation learning technique. This approach can capture local structural motifs but is highly sensitive to the choice of clustering parameters, and it may struggle to adequately represent more complex ring structures, which are critical in many chemical applications.
In this work, we introduce a functional group detection algorithm that uses rule-based methods to accurately identify 101 common functional groups as well as all ring-containing functional groups. Given a molecule, we can assign each atom to a functional group, ensuring that every atom is associated with a functional unit. This comprehensive assignment enables more nuanced tokenization and fragmentation by accounting for the chemical context of each atom, ultimately producing more meaningful and chemically consistent representations for downstream tasks.
2.2 Contrastive Learning-Based Molecular Representations
Contrastive learning, a self-supervised learning technique, aims to learn representations by maximizing agreement between positive pairs while distinguishing them from negative samples. In the context of molecular representation, Wang et al. (2022b) applies this technique by augmenting each molecular graph to create slightly different versions, which are then treated as negative examples to enhance the learning process. In Pinheiro et al. (2022), the authors use a molecular graph encoder and a SMILES encoder to encode both molecular graphs and SMILES, employing contrastive loss to maximize the agreement between these embeddings. This method enriches SMILES with topology information and the molecular graph with sequence context, making the final embeddings of molecules more robust. In Zhang et al. (2020), the authors minimize the distance between the representation of a molecule and the representations of its constituent substructures. Collectively, these works highlight the versatility of contrastive learning in unifying diverse molecular representations, leading to improved downstream performance in molecular tasks. Our method extends this approach by using contrastive learning to maximize the agreement between FG-enhanced SMILES representations and FG graphs, thereby enhancing the alignment between sequence-based and graph-based representations, ensuring a more comprehensive and chemically informed embedding of molecules.
@@ 3 Methodology
In this section, we present our methodologies for enhancing molecular representation learning. Section 3.1 introduces FG-aware tokenization and fragmentation, which injects detailed chemical context into both sequence and graph representations. Section 3.2 delves into the masked atom prediction task, a self-supervised technique that enables the model to learn atom-level representations. In Section 3.3, we describe our method for capturing the core molecular structure. Finally, Section 3.4 presents contrastive learning, which aligns FG-enhanced SMILES strings with FG graph embeddings to achieve a comprehensive, unified molecular representation.
@@ 3.1 Functional group-Aware Tokenization and Fragmentation
We propose an FG-aware tokenization and fragmentation method that embeds detailed functional group information into molecular representations, tailoring it for both SMILES and graph-based models. This method defines a set of functional groups and employs an algorithm to detect them within the molecular graph. The algorithm traverses the graph, evaluating each atom based on criteria such as atom type, neighboring atom types and corresponding bonds, number of neighbors, atom charge, and bonded hydrogen atoms to identify the functional group. This approach ensures precise detection of functional groups within molecular graphs. For instance, a carbon atom with a charge of 00 and three neighbors, including a double-bonded oxygen atom, is classified as part of a ketone group (RCOR’), with the oxygen also contributing to the group. Appendix B.1 provides a full list of non-ring-containing functional groups.
After identifying conventional functional groups, such as hydroxyl (-OH) and carboxyl (-COOH), the molecule is further analyzed to detect rings and fused ring systems, which are also treated as functional groups. These ring systems are highly diverse and make up a significant portion of all functional groups. For atoms that cannot be assigned to any predefined functional group—typically rare atoms like Ag or Fe—their chemical symbols are used to represent the functional group. This ensures that all atoms are included in the functional group representation, even if they do not fit into standard categories. Once all functional groups are identified, the molecule is segmented at bonds connecting these groups, preserving the functional groups’ integrity during fragmentation. The results of the FG detection algorithm are utilized for two distinct processes:
  • FG-aware tokenization: The molecular graph, where each node (atom) is assigned to a specific functional group, is converted back into a SMILES string that incorporates functional group information. For example, a ketone-containing group originally depicted in SMILES as *C(=O)* is transformed into an FG-enhanced SMILES string like *C_ketone(=O_ketone)*. This FG-enhanced SMILES embeds additional chemical context directly into the molecular representation while remaining fully compliant with traditional SMILES rules (if the FG information is removed, the FG-enhanced SMILES reverts to its standard SMILES form).
@@ 3.2 Atom-level Representation Learning
We employ a masked language model architecture that takes FG-enhanced SMILES as input, using atom-level tokenization and leveraging masked atom prediction as a self-supervised task to train the model. Specifically, we adopt the BERT (Devlin, 2018) with a masked language modeling loss:
$\mathcal{L}_{\text{MLM}}=-\sum_{i\in\mathcal{M}}\log P_{\theta}(x_{i}\mid x_{\setminus i}),$
@@ 3.3 Molecular Structure Learning
The FG knowledge graph embedding is learned by the ComplEx model (Trouillon et al., 2016) to obtain node embeddings. ComplEx is a matrix factorization model specifically designed to learn embeddings from multi-relational data. It is particularly effective at capturing complex, asymmetric relationships and handling one-to-many relations in knowledge graphs, making it well-suited for modeling intricate relational structures. Specifically, each element in a triple $(\mathbf{h},\mathbf{r},\mathbf{t})$ — where $\mathbf{h}$ is the head entity, $\mathbf{r}$ is the relation, and $\mathbf{t}$ is the tail entity — is represented as a complex vector. The score for a given triple $(\mathbf{h},\mathbf{r},\mathbf{t})$ is calculated as:
$f(\mathbf{h},\mathbf{r},\mathbf{t})=\text{Re}\left(\mathbf{h}^{T}\mathbf{r}\cdot\mathbf{t}\right)$
ComplEx employs a margin-based ranking loss function defined as:
$\mathcal{L}_{\text{Graph}}=\sum_{(\mathbf{h},\mathbf{r},\mathbf{t})\in E^{+}}\sum_{(\mathbf{h^{\prime}},\mathbf{r},\mathbf{t^{\prime}})\in E^{-}}\max\left(0,\gamma+f(\mathbf{h^{\prime}},\mathbf{r},\mathbf{t^{\prime}})-f(\mathbf{h},\mathbf{r},\mathbf{t})\right)$
@@
where $\mathbf{h}_{i}^{\prime}$ is the updated embedding for node $i$, computed by averaging the embeddings $\mathbf{h}_{j}$ of neighboring nodes $\mathcal{N}(i)$, applying the weight matrix $\mathbf{W}$, and passing through ReLU activation function. The score estimates the probability of connections between nodes is computed with a multi-layer perceptron (MLP) and a sigmoid function:
$p_{ij}=\sigma(\text{MLP}(\mathbf{h}_{i}\oplus\mathbf{h}_{j}))$
@@
Figure 5(b) and Figure 9 (Appendix D.3) demonstrates the capability of our molecular structure representation model, showing that, akin to word pair analogy tasks in NLP, replacing one functional group in a molecule with another produces parallel results across different molecules. This demonstrates the model’s ability to effectively capture and preserve chemical analogies, highlighting its robustness in learning and representing molecular structures.

### v3

@@ Abstract
We introduce Functional Group-Aware Representations for Small Molecules (FARM), a novel foundation model designed to bridge the gap between SMILES, natural language, and molecular graphs. The key innovation of FARM lies in its functional group-aware tokenization, which directly incorporates functional group information into SMILES, enriching SMILES with detailed chemical context. For example, instead of using “O” to represent all oxygen atoms, we use specific tokens like “O_ketone” and “O_hydroxyl” to differentiate oxygen atoms belonging to distinct functional groups. This tokenization expands the chemical lexicon, effectively bridging the gap between SMILES and natural language in terms of vocabulary size, ultimately enhancing the model’s ability to predict molecular properties. FARM also represents molecules from two perspectives: by (1) using masked language modeling to capture atom-level features and (2) employing graph neural networks to encode the whole molecule topology. FARM leverages contrastive learning to aligns these two views of representations into a unified molecular embedding. We rigorously evaluate FARM on the MoleculeNet dataset, where it achieves state-of-the-art performance on 11 out of 13 tasks. These results highlight FARM’s potential to improve molecular representation learning and demonstrate its strong transfer learning capabilities, paving the way for promising applications in drug discovery and pharmaceutical research. The code is available at: https://github.com/thaonguyen217/farm_molecular_representation.
@@ 1 Introduction
Artificial intelligence (AI) has emerged as a transformative tool in accelerating scientific discovery, particularly in drug development. It is increasingly employed for tasks such as molecular property prediction, drug-target interaction prediction, and quantitative structure-activity relationship (QSAR) modeling (Chen et al., 2016; Wen et al., 2017; Shen and Nicolaou, 2019; Walters and Barzilay, 2020; Achary, 2020; Wang et al., 2022a; Edwards et al., 2022; Zhang et al., 2023a; Nguyen et al., 2024a; Edwards et al., 2024b, a). However, one of the central challenges in this field is the scarcity of large labeled datasets required for traditional supervised learning methods. This has shifted the focus towards self-supervised pre-trained models that can extract meaningful patterns from vast amounts of unlabeled molecular data (Shen and Nicolaou, 2019). As a result, the development of robust foundation models for molecular representations is now more critical than ever.
Despite significant advancements in other domains, such as natural language processing (NLP) and computer vision, there is still no dominant foundation model tailored to molecular representation in drug discovery (Zhang et al., 2023b). This paper begins to address this pressing gap by introducing an innovative approach that leverages functional group (FG)-aware tokenization in the context of both sequence-based and graph-based molecular representations. In the literature, terms such as “motifs,” “fragments,” “substructures,” and “building blocks” are often used interchangeably with “functional groups.” However, in this paper, we use the term “functional groups” to represent a subset of these molecular concepts, with a more rigorous definition grounded in chemical principles. A functional group refers to a chemically meaningful portion of a molecule that significantly influences its properties and behavior. This can include simple functional groups, such as hydroxyl (-OH), as well as more complex molecular substructures, such as ring systems, which serve distinct functional roles within the molecule. Functional groups (FGs) play a crucial role in determining a molecule’s properties. For instance, salicylic acid and aspirin share the same core structure but differ by just one functional group—a hydroxyl group in salicylic acid is replaced by an acetyl group in aspirin. This seemingly minor change results in substantial differences in their chemical properties and biological activities. The acetyl group in aspirin enables it to irreversibly inhibit cyclooxygenase (COX) enzymes, providing it with blood-thinning properties that make it effective in preventing heart attacks and strokes—functions that salicylic acid does not exhibit.
The innovation here lies in expanding tokenization granularity for SMILES and reducing tokenization granularity for molecular graphs in a way that explicitly interfaces with key drivers of molecular function. This approach enables molecular embeddings to more effectively support models in understanding and predicting molecular functions. For example, instead of representing all oxygen atoms with the same token (e.g., “O”), we use more specific tokens such as “O_ketone” and “O_hydroxyl” to distinguish oxygen atoms that belong to different functional groups in the SMILES string. This approach addresses a major limitation of sequence-based models, which typically focus only on individual atom types while neglecting the higher-level functional groups crucial for accurate molecular understanding. FG-aware tokenization enriches the SMILES representation with chemically relevant context, bridging the gap between the expansive vocabulary of natural language—the language for which Transformers were originally designed—and the limited chemical lexicon of the SMILES language, thereby reducing negative transfer during learning.
Conversely, instead of representing a ketone group (C=O) using two separate nodes in a graph, we represent it as a single node corresponding to the functional group as a whole. This abstraction yields a high-level “skeleton” of the molecule, where each node denotes a chemically meaningful functional unit rather than individual atoms. To capture interactions between these units, we construct a functional group (FG) knowledge graph that encodes chemically informed embeddings for each functional group based on their properties and co-occurrence patterns. These embeddings are then leveraged in a link prediction task to learn structural representations of molecules—allowing the model to infer both functional group similarity and likely connectivity. This FG graph captures structural relationships that are often implicit or obscured in SMILES representations. As a result, the functional group graph provides a complementary perspective to the FG-enhanced SMILES, encoding topological and structural context, while the enhanced SMILES preserves fine-grained sequence-level and chemical detail. Contrastive learning will then be used to align these two views, encouraging the model to learn representations that are sensitive to chemically meaningful structures. Figure 1 and Figure 2a present an overview of our molecular representation learning model.
Figure 1: FARM’s molecular representation learning framework. Given a molecule as input, we apply FG-aware tokenization to its SMILES representation and FG-based fragmentation to its molecular graph. Each resulting representation is then processed through a self-supervised learning objective. Finally, the two views are aligned using contrastive learning to make the final embedding robust by capturing both atom-level information from the SMILES and structural information from the graph.
@@
  • Robustness in downstream tasks: FARM demonstrates strong transfer learning capabilities - the core goal of pretrained models - outperforming other methods in 11 out of 13 tasks from the MoleculeNet benchmark.
2 Related Work
Functional group-Aware Molecular Representations   In recent years, there has been growing recognition that incorporating functional group (FG) information into molecular representations can significantly enhance model performance in downstream tasks. Approaches in this domain can be broadly classified into two categories: those leveraging language models (LMs) (Li et al., 2023; Xia et al., 2023) and those utilizing graph neural networks (GNNs) (Zhang et al., 2020, 2021; Yu and Gao, 2022; Yang et al., 2022; Wang et al., 2023; Wu et al., 2023; Han et al., 2023; Fang et al., 2023). Within these categories, methods either employ rule-based functional group detection, relying on predefined chemical rules to identify FGs, or adopt unsupervised strategies that infer substructures or motifs from the data. Regardless of the approach, these models consistently demonstrate that enriching molecular representations with functional group information leads to improved performance across a wide range of molecular property prediction tasks (Fang et al., 2023; Han et al., 2023; Wang et al., 2023). This underlines the importance of functional group awareness in achieving accurate and generalizable molecular representations.
In studies such as Zhang et al. (2021); Han et al. (2023); Chen et al. (2024); Yang et al. (2022), the BRICS algorithm (Degen et al., 2008) is employed to fragment molecules, with some extending the approach through additional rules to achieve finer-grained segmentation. However, this method focuses on general reaction-based bond breaking, rather than targeting functional group-specific characteristics. Additionally, while BRICS groups atoms into fragments, it does not explicitly identify or label functional groups, limiting its utility in tasks that require functional groups-specific information. Other works (Li et al., 2023; Chen et al., 2024) leverage RDKit (Landrum, 2010) for functional group detection. While RDKit is effective in identifying common functional groups, its capabilities are limited to well-known groups and do not extend to detecting more complex functional groups, such as ring systems that often dominate molecular datasets and play a critical role in molecular function. In this work, we present a rule-based functional group detection algorithm that accurately identifies 101 common groups and all ring-containing ones. Unlike frequent subgraph mining or motif-based methods, our approach is explicitly grounded in chemical principles, ensuring reliable and chemically meaningful detection aligned with expert definitions.
Contrastive Learning-Based Molecular Representations   Contrastive learning, a self-supervised learning technique, aims to learn representations by maximizing agreement between positive pairs while distinguishing them from negative samples. In the context of molecular representation, Wang et al. (2022b) applies this technique by augmenting each molecular graph to create slightly different versions, which are then treated as negative examples to enhance the learning process. In Pinheiro et al. (2022), the authors use a molecular graph encoder and a SMILES encoder to encode both molecular graphs and SMILES, employing contrastive loss to maximize the agreement between these embeddings. This method enriches SMILES with topology information and the molecular graph with sequence context, making the final embeddings of molecules more robust. In Zhang et al. (2020), the authors minimize the distance between the representation of a molecule and the representations of its constituent substructures. Collectively, these works highlight the versatility of contrastive learning in unifying diverse molecular representations, leading to improved downstream performance in molecular tasks. Our method extends this approach by using contrastive learning to maximize the agreement between FG-enhanced SMILES representations and FG graphs, thereby enhancing the alignment between sequence-based and graph-based representations, ensuring a more comprehensive and chemically informed embedding.
@@ 3 Methodology
We present our methodologies for enhancing molecular representation learning. Section 3.1 introduces FG-aware tokenization and fragmentation, which injects detailed chemical context into both sequence and graph representations. Section 3.2 delves into the masked atom prediction task, a self-supervised technique that enables the model to learn atom-level representations. In Section 3.3, we describe our method for capturing the core molecular structure. Finally, Section 3.4 presents contrastive learning, which aligns FG-enhanced SMILES strings with FG graph embeddings to achieve a comprehensive, unified molecular representation.
@@ 3.1 Functional Group-Aware Tokenization and Fragmentation
We propose an FG-aware tokenization and fragmentation method that embeds detailed functional group information into molecular representations, tailoring it for both SMILES and graph-based models. This method defines a set of functional groups and employs an algorithm to detect them within the molecular graph. The algorithm traverses the graph, evaluating each atom based on criteria such as atom type, neighboring atom types and corresponding bonds, number of neighbors, atom charge, and bonded hydrogen atoms to identify the functional group. This approach ensures precise detection of functional groups within molecular graphs. For instance, a carbon atom with a charge of 0 and three neighbors, including a double-bonded oxygen atom, is classified as part of a ketone group (RCOR’), with the oxygen also contributing to the group. Additionally, we address cases where a functional group may be a subset of another. The algorithm first checks for the presence of the larger functional group. If it’s not identified, the algorithm then checks for smaller functional groups, ensuring correct identification even in complex structures. Appendix B.1 provides a full list of 101 non-ring-containing functional groups.
The results of the FG detection algorithm are utilized for two distinct processes:
  • FG-aware tokenization: The molecular graph, where each node (atom) is assigned to a specific functional group, is converted back into a SMILES string that incorporates functional group information. For example, a ketone-containing group originally depicted in SMILES as *C(=O)* is transformed into an FG-enhanced SMILES string like *C_ketone(=O_ketone)*. This FG-enhanced SMILES embeds additional chemical context directly into the molecular representation while remaining fully compliant with traditional SMILES rules (if the FG information is removed, the FG-enhanced SMILES reverts to its standard SMILES form). Similar to how words like “bank” have multiple meanings depending on context (e.g., river bank vs. financial bank), our method captures fine-grained semantic distinctions by representing functional groups with more precise tokens—for example, distinguishing ketone oxygen from hydroxyl oxygen.
@@ 3.2 Atom-Level Representation Learning
We employ a masked language model architecture that takes FG-enhanced SMILES as input, using atom-level tokenization and leveraging masked atom prediction as a self-supervised task to train the model. Specifically, we adopt the BERT (Devlin, 2018) with the following loss:
$\mathcal{L}_{\text{MLM}}=-\sum_{i\in\mathcal{M}}\log P_{\theta}(x_{i}\mid x_{\setminus i}),\vskip-1.99997pt$
@@ 3.3 Molecular Structure Learning
The FG knowledge graph embedding is learned by the ComplEx model (Trouillon et al., 2016) to obtain node embeddings. ComplEx is a matrix factorization model specifically designed to learn embeddings from multi-relational data. It is particularly effective at capturing complex, asymmetric relationships and handling one-to-many relations in knowledge graphs, making it well-suited for modeling intricate relational structures. Specifically, each element in a triple $(\mathbf{h},\mathbf{r},\mathbf{t})$ — where $\mathbf{h}$ is the head entity, $\mathbf{r}$ is the relation, and $\mathbf{t}$ is the tail entity — is represented as a complex vector. The score for a given triple $(\mathbf{h},\mathbf{r},\mathbf{t})$ is calculated as: $f(\mathbf{h},\mathbf{r},\mathbf{t})=\text{Re}\left(\mathbf{h}^{T}\mathbf{r}\cdot\mathbf{t}\right)$. ComplEx employs a margin-based ranking loss function defined as:
$\mathcal{L}_{\text{Graph}}=\sum_{\begin{subarray}{c}(\mathbf{h},\mathbf{r},\mathbf{t})\\ \in E^{+}\end{subarray}}\sum_{\begin{subarray}{c}(\mathbf{h^{\prime}},\mathbf{r},\mathbf{t^{\prime}})\\ \in E^{-}\end{subarray}}\max\left(0,\gamma+f(\mathbf{h^{\prime}},\mathbf{r},\mathbf{t^{\prime}})-f(\mathbf{h},\mathbf{r},\mathbf{t})\right)$ (1)
@@
where $\mathbf{h}_{i}^{\prime}$ is the updated embedding for node $i$, computed by averaging the embeddings $\mathbf{h}_{j}$ of neighboring nodes $\mathcal{N}(i)$, applying the weight matrix $\mathbf{W}$, and passing through ReLU activation function. The score estimates the probability of connections between nodes is computed with a multi-layer perceptron (MLP) and a sigmoid function: $p_{ij}=\sigma(\text{MLP}(\mathbf{h}_{i}\oplus\mathbf{h}_{j}))$.
@@
Figure 4(b) and Figure 8 (Appendix D.3) demonstrates the capability of our molecular structure representation model, showing that, akin to word pair analogy tasks in NLP, replacing one functional group in a molecule with another (in this case, replacing -OH with -COOH) produces parallel results across different molecules. This demonstrates the model’s ability to effectively capture and preserve chemical analogies, highlighting its robustness in learning and representing molecular structures.